​(Operating leverage​) The C. M. Quarles Distributing Company manufactures an assortment of cold air intake systems for​ high-performance engines. The average selling price for the various units is ​$650. The associated variable cost is ​$250 per unit. Fixed costs for the firm average $ 170,000 annually.

a. What is the​ break-even point in units for the​ company?

b. What is the dollar sales volume the firm must achieve to reach the​ break-even point?

c. What is the degree of operating leverage for a production and sales level of 6,000 units for the​ firm? (Calculate to three decimal​ places.)

d. What will be the projected effect on earnings before interest and taxes if the​ firm's sales level should increase by 35 percent from the volume noted in part c​?
Answer #1

a. What is the​ break-even point in units for the​ company?
=170000/(650-250)
=425

b. What is the dollar sales volume the firm must achieve to reach the​ break-even point?
=425*650
=276250

c. What is the degree of operating leverage for a production and sales level of 6,000 units for the​ firm? (Calculate to three decimal​ places.)
=6000*(650-250)/(6000*(650-250)-170000)
=1.076233184

d. What will be the projected effect on earnings before interest and taxes if the​ firm's sales level should increase by 35 percent from the volume noted in part c​?
% increase in EBIT =1.076233184*35%
=37.67%